OpenStudy (samigupta8):

A carnot engine having efficiency 40% used as a refrigerator. If work done by compressor of gas is 100J.Find heat absorbed from the body having lower temperature.

OpenStudy (samigupta8):

I got the ans as 150J. And it is wrong

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (big_mensa961):

@QueenFromADayOfRace @nubeer

OpenStudy (irishboy123):

http://www.ohio.edu/people/piccard/phys202/carnot/carnot.html if that link is correct then you are looking at \(0.4 = \dfrac{Q_e}{100 - Q_e}\) scroll down to fridge

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

3rd Last post i made

OpenStudy (samigupta8):

Further beta is q2/w Q2 is heat absorbed from the sink W is work done on the fridge by external source

OpenStudy (samigupta8):

Beta=3/2 And if i substitute it into Beta=q2/w

OpenStudy (samigupta8):

We get q2=150J

OpenStudy (samigupta8):

@irishboy123 pls help

OpenStudy (samigupta8):

@pawanyadav